The diagonal of a rectangle is 13 cm. How long is its width if its length is 10 cm?
Answer by stanbon(75887)

The diagonal of a rectangle is 13 cm. How long
is its width if its length is 10 cm?
---------------------
Draw the picture:
You have a right triangle with hypotenuse = 13 and one side is 10
=====
Equation:
Let the length be "x":
13^2 = 10^2 + x^2
-----
169 = 100 + x^2
------
x^2 = 69
----
length = sqrt(69)
=======================
Cheers,
Stan H.
======================